# An Explicit M1 Radiation-hydrodynamics Scheme for Three-dimensional Protostellar Evolution

### v1

@@ Appendix A Lorentz transformations
which are obtained by retaining terms up to first order in $\beta$ in the Lorentz transformations (see Equations 91.10 and 91.11 in D. Mihalas & B. W. Mihalas 1984). In Equation (A2), the second term on the right-hand side represents the advection of radiation enthalpy by the moving gas.
@@ Appendix B Mixed frame RHD Equations at Leading Order
In this appendix, we derive the set of equations that should be solved when the gas velocity is much smaller than the speed of light, $\beta\ll 1$. A key point to note is that we aim to perform the simulation resolving regions inside protostars, where the density is high and the medium is extremely optically thick ($\tau\gg 1$). In this case, terms proportional to $\beta$ are not necessarily negligible since the product $\tau\beta$ can be very large, which is the so-called dynamic diffusion regime. Therefore we must take special care in the derivation of the RHD equations. A detailed analysis of the magnitude of the terms appearing in the RHD equations under the assumption of $\beta\ll 1$ has been carried out by M. R. Krumholz et al. (2007b) and M. A. Skinner & E. C. Ostriker (2013). They examine the relative magnitudes of each term in the equations across various regimes, including the optically thin limit ($\tau\ll 1$), the static diffusion limit ($\tau\gg 1$ and $\tau\beta\ll 1$), and the dynamic diffusion limit ($\tau\gg 1$ and $\tau\beta\gg 1$). In this paper, based on their analysis, we derive the RHD equations by retaining only the leading-order terms.
First, by retaining the terms up to the second order in $\beta$ in Equations (54b) and (54d) of D. Mihalas & L. Auer (2001), we obtain the radiation four-force $(G^{0},\bm{G})$ density as
@@
Considering the optically thick regime and replacing $\kappa_{0E}$ and $\chi_{0F}$ (Equations 9 and 10) with $\kappa_{0\mathrm{P}}$ and $\chi_{0\mathrm{R}}$ (Equations 8 and 12) make these equations equivalent to Equations (24) and (25) in M. R. Krumholz et al. (2007b).
@@
In Table 1, we present the scaling of each term in Equation (B1) and (B2). Our scaling analysis yields results identical to those in Table 1 of M. R. Krumholz et al. (2007b), except that we adopt mean opacities for general spectra instead of those for thermal spectra considered in M. R. Krumholz et al. (2007b). For $G^{0}$ in the dynamic diffusion regime, the first, second, and third terms on the right-hand side of Equation (B1) can all become dominant. However, it is not necessary to solve the second and third terms if the goal is to obtain the radiation energy density $E_{\mathrm{rad}}$ to leading order. As shown in the first row of Table 1, the quantity $\rho(\kappa_{0E}E_{\mathrm{rad}}-\kappa_{0\mathrm{P}}aT^{4})(l/E_{\mathrm{rad,0}})$ scales as $\mathcal{O}(\beta^{2}\tau)$. Assuming $\tau\gg 1$ and $\kappa_{0E}=\kappa_{0\mathrm{P}}$, this implies that $E_{\mathrm{rad}}$ deviates from $aT^{4}$ only by an amount of order $\mathcal{O}(\beta^{2}E_{\mathrm{rad,0}})$. While resolving this small deviation requires inclusion of the second and third terms, such accuracy is unnecessary if one is concerned only with the leading-order behavior. Therefore, the second and third terms can be neglected in that case. On the other hand, for the momentum source term $\bm{G}$ given by Equation (B2), the dominant contributions arise from the first and second terms on the right-hand side. Thus, we need to retain both of these terms.
@@ Appendix C Equations for Photon Number Density
We estimate $\chi^{\prime}_{0F}$ in the same manner as $\chi_{0F}$ in Section 4.3, obtaining
@@ Appendix D Impact of a Non-conservative Self-gravity Scheme on Protostellar Evolution
Figure 8: Comparison of spherical collapse simulations using the energy-conservative self-gravity scheme of P. D. Mullen et al. (2021) and a conventional non-conservative scheme. The left panel shows the time evolution of the maximum density $\rho_{\mathrm{max}}$ in the computational domain, which corresponds to the central density of the protostar after its formation. The right panel presents the total energy $\mathcal{E}$ that should be conserved during the calculation, with its definition given in the text. Since $\mathcal{E}$ takes negative values due to the dominance of gravitational energy, we plot -$\mathcal{E}$ in the figure. The blue and orange lines indicate the results obtained with the energy-conservative and non-conservative schemes, respectively. The moment when the central density reaches $2.4\times 10^{-7}~\mathrm{g~cm}^{-3}$ is defined as $t=0$.
As described in Section 5.2, we adopt the self-gravity scheme proposed by P. D. Mullen et al. (2021), which ensures total energy conservation with higher accuracy than the conventional approach and preserves the curl-free nature of the gravitational field, $\nabla\times\bm{g}=0$. Unlike conventional self-gravity updates, which treat gravity simply as source terms, their method evaluates the source terms in a way that is mathematically equivalent to flux divergences. In this section, we compare the protostellar evolution obtained with this energy-conservative scheme to that from a conventional non-conservative one. To this end, we simulate the formation and evolution of a protostar resulting from the collapse of a spherical cloud, incorporating primordial gas chemistry and modeling radiation with two frequency bins, following the approach described in Section 7. As the initial condition, we adopt a critical isothermal Bonnor–Ebert sphere with a central density of $2.3\times 10^{-15}~\mathrm{g~cm}^{-3}$ and a temperature of $4.8\times 10^{3}$ K (W. B. Bonnor, 1956). These values are taken from the evolutionary path of a collapsing primordial cloud under strong far-ultraviolet radiation, as given by the one-zone calculations of K. Omukai (2001). To induce collapse, the density profile is enhanced by a factor of $1.6$. The radius of the Bonnor–Ebert sphere is $5.4\times 10^{3}~\mathrm{au}$, while the length of the entire computational domain is $1.8\times 10^{4}~\mathrm{au}$. The gas is kept static outside the Bonnor–Ebert sphere, whereas radiation is solved throughout the domain and allowed to escape freely through the outer boundary. The computational domain is adaptively refined to resolve the local Jeans length with at least 8 cells. We then follow the gravitational collapse and investigate the subsequent evolution of the protostar formed at the cloud center.
@@
where the superscript $n$ denotes the time step index, $\Delta t^{n}$ is the value of the $n$-th time step, $d\bm{S}$ is the area element, and the integration is performed over the boundary $\partial V$ of the computational domain. Although the total energy $\mathcal{E}_{\mathrm{tot}}$ is not strictly conserved because the gas outside the Bonnor–Ebert sphere is artificially kept static by the boundary condition, the central high-density region dominates the energy budget. Therefore, the influence of the boundary treatment is negligible, and the total energy is expected to remain constant. In the present setting, the gravitational energy dominates, and thus $\mathcal{E}_{\mathrm{tot}}$ takes negative values. As seen in the right panel of Figure 8, in the conservative scheme, $\mathcal{E}_{\mathrm{tot}}$ remains nearly constant with the relative difference between $t=0$ and $t=1$ being only $10^{-4}$. In contrast, in the non-conservative scheme, $\mathcal{E}_{\mathrm{tot}}$ is larger than in the conservative case throughout the evolution, and after protostar formation it gradually increases with time, i.e., its absolute value decreases, indicating a net gain of energy. Such an artificial increase in energy and the associated density decrease has also been reported by P. D. Mullen et al. (2021) in hydrostatic systems without accretion. These results highlight the importance of employing a conservative self-gravity scheme to accurately follow the evolution of self-gravitating systems such as protostars.

### v2

@@ Appendix A Lorentz transformations
which are obtained by retaining terms up to first order in $\beta$ in the Lorentz transformations (see Equations 91.10 and 91.11 in Mihalas and Mihalas 1984). In Equation (A2), the second term on the right-hand side represents the advection of radiation enthalpy by the moving gas.
@@ Appendix B Mixed frame RHD Equations at Leading Order
In this appendix, we derive the set of equations that should be solved when the gas velocity is much smaller than the speed of light, $\beta\ll 1$. A key point to note is that we aim to perform the simulation resolving regions inside protostars, where the density is high and the medium is extremely optically thick ($\tau\gg 1$). In this case, terms proportional to $\beta$ are not necessarily negligible since the product $\tau\beta$ can be very large, which is the so-called dynamic diffusion regime. Therefore we must take special care in the derivation of the RHD equations. A detailed analysis of the magnitude of the terms appearing in the RHD equations under the assumption of $\beta\ll 1$ has been carried out by Krumholz et al. (2007a) and Skinner and Ostriker (2013). They examine the relative magnitudes of each term in the equations across various regimes, including the optically thin limit ($\tau\ll 1$), the static diffusion limit ($\tau\gg 1$ and $\tau\beta\ll 1$), and the dynamic diffusion limit ($\tau\gg 1$ and $\tau\beta\gg 1$). In this paper, based on their analysis, we derive the RHD equations by retaining only the leading-order terms.
First, by retaining the terms up to the second order in $\beta$ in Equations (54b) and (54d) of Mihalas and Auer (2001), we obtain the radiation four-force $(G^{0},\bm{G})$ density as
@@
Considering the optically thick regime and replacing $\kappa_{0E}$ and $\chi_{0F}$ (Equations 9 and 10) with $\kappa_{0\mathrm{P}}$ and $\chi_{0\mathrm{R}}$ (Equations 8 and 12) make these equations equivalent to Equations (24) and (25) in Krumholz et al. (2007a).
@@
In Table 6, we present the scaling of each term in Equation (B1) and (B2). Our scaling analysis yields results identical to those in Table 1 of Krumholz et al. (2007a), except that we adopt mean opacities for general spectra instead of those for thermal spectra considered in Krumholz et al. (2007a). For $G^{0}$ in the dynamic diffusion regime, the first, second, and third terms on the right-hand side of Equation (B1) can all become dominant. However, it is not necessary to solve the second and third terms if the goal is to obtain the radiation energy density $E_{\mathrm{rad}}$ to leading order. As shown in the first row of Table 6, the quantity $\rho(\kappa_{0E}E_{\mathrm{rad}}-\kappa_{0\mathrm{P}}aT^{4})(l/E_{\mathrm{rad,0}})$ scales as $\mathcal{O}(\beta^{2}\tau)$. Assuming $\tau\gg 1$ and $\kappa_{0E}=\kappa_{0\mathrm{P}}$, this implies that $E_{\mathrm{rad}}$ deviates from $aT^{4}$ only by an amount of order $\mathcal{O}(\beta^{2}E_{\mathrm{rad,0}})$. While resolving this small deviation requires inclusion of the second and third terms, such accuracy is unnecessary if one is concerned only with the leading-order behavior. Therefore, the second and third terms can be neglected in that case. On the other hand, for the momentum source term $\bm{G}$ given by Equation (B2), the dominant contributions arise from the first and second terms on the right-hand side. Thus, we need to retain both of these terms.
@@ Appendix C Equations for Photon Number Density
We estimate $\chi^{\prime}_{0F}$ in the same manner as $\chi_{0F}$ in Section IV.3, obtaining
@@ Appendix D Impact of a Non-conservative Self-gravity Scheme on Protostellar Evolution
Figure 9: Comparison of spherical collapse simulations using the energy-conservative self-gravity scheme of Mullen et al. (2021) and a conventional non-conservative scheme. The left panel shows the time evolution of the maximum density $\rho_{\mathrm{max}}$ in the computational domain, which corresponds to the central density of the protostar after its formation. The right panel presents the total energy $\mathcal{E}$ that should be conserved during the calculation, with its definition given in the text. Since $\mathcal{E}$ takes negative values due to the dominance of gravitational energy, we plot -$\mathcal{E}$ in the figure. The blue and orange lines indicate the results obtained with the energy-conservative and non-conservative schemes, respectively. The moment when the central density reaches $2.4\times 10^{-7}~\mathrm{g~cm}^{-3}$ is defined as $t=0$.
As described in Section V.2, we adopt the self-gravity scheme proposed by Mullen et al. (2021), which ensures total energy conservation with higher accuracy than the conventional approach and preserves the curl-free nature of the gravitational field, $\nabla\times\bm{g}=0$. Unlike conventional self-gravity updates, which treat gravity simply as source terms, their method evaluates the source terms in a way that is mathematically equivalent to flux divergences. In this section, we compare the protostellar evolution obtained with this energy-conservative scheme to that from a conventional non-conservative one. To this end, we simulate the formation and evolution of a protostar resulting from the collapse of a spherical cloud, incorporating primordial gas chemistry and modeling radiation with two frequency bins, following the approach described in Section VII. As the initial condition, we adopt a critical isothermal Bonnor–Ebert sphere with a central density of $2.3\times 10^{-15}~\mathrm{g~cm}^{-3}$ and a temperature of $4.8\times 10^{3}$ K (Bonnor, 1956). These values are taken from the evolutionary path of a collapsing primordial cloud under strong far-ultraviolet radiation, as given by the one-zone calculations of Omukai (2001). To induce collapse, the density profile is enhanced by a factor of $1.6$. The radius of the Bonnor–Ebert sphere is $5.4\times 10^{3}~\mathrm{au}$, while the length of the entire computational domain is $1.8\times 10^{4}~\mathrm{au}$. The gas is kept static outside the Bonnor–Ebert sphere, whereas radiation is solved throughout the domain and allowed to escape freely through the outer boundary. The computational domain is adaptively refined to resolve the local Jeans length with at least 8 cells. We then follow the gravitational collapse and investigate the subsequent evolution of the protostar formed at the cloud center.
@@
where the superscript $n$ denotes the time step index, $\Delta t^{n}$ is the value of the $n$-th time step, $d\bm{S}$ is the area element, and the integration is performed over the boundary $\partial V$ of the computational domain. Although the total energy $\mathcal{E}_{\mathrm{tot}}$ is not strictly conserved because the gas outside the Bonnor–Ebert sphere is artificially kept static by the boundary condition, the central high-density region dominates the energy budget. Therefore, the influence of the boundary treatment is negligible, and the total energy is expected to remain constant. In the present setting, the gravitational energy dominates, and thus $\mathcal{E}_{\mathrm{tot}}$ takes negative values. As seen in the right panel of Figure 9, in the conservative scheme, $\mathcal{E}_{\mathrm{tot}}$ remains nearly constant with the relative difference between $t=0$ and $t=1$ being only $10^{-4}$. In contrast, in the non-conservative scheme, $\mathcal{E}_{\mathrm{tot}}$ is larger than in the conservative case throughout the evolution, and after protostar formation it gradually increases with time, i.e., its absolute value decreases, indicating a net gain of energy. Such an artificial increase in energy and the associated density decrease has also been reported by Mullen et al. (2021) in hydrostatic systems without accretion. These results highlight the importance of employing a conservative self-gravity scheme to accurately follow the evolution of self-gravitating systems such as protostars.